OpenStudy (anonymous):

Please help on finding limits? 1) Find: lim (x-->infinity) 9x^4+3/4x^4 A) 3/4 B) 4/9 C) 9/4 D) 0 Please explain step by step. I'm really interested in learning how to solve this. Thank you <3 2) Use a graphing calculator to find the value of the limit (use radians with trigonometric functions) lim (x-->4) -2sqrt(x)+6/9-x A) 0.9 B) 0.1 C) -0.4 (this is the one I believe is correct from what I tried to do) D) 0.4 Please explain step by step also. I have a TI-83 Plus.

OpenStudy (solomonzelman):

1) L'H'S

OpenStudy (zarkon):

\[\frac{9x^4+3}{4x^4}=\frac{9x^4}{4x^4}+\frac{3}{4x^4}=\frac{9}{4}+\frac{3}{4x^4}\]